# BeFA: A General Behavior-driven Feature Adapter for Multimedia Recommendation

### v1

@@ 4.1. Problem Formulation
ImageBind Avg Imporve 15.54% 9.09% 17.99% 12.53% 5.18% 3.77% 4.83% 4.12% 15.24% 13.87% 19.56% 17.92%
4.2. Behavior-driven Feature Adapter
@@
Due to the shortcomings of pre-trained feature encoders, their extracted content features contain a significant amount of irrelevant and erroneous information. To better utilize modality information,we propose Behaviour-driven Feature Adapter(BeFA) for adapting content features. Firstly, we decouple the original item content features $\mathrm{E_{i,m}}$ into features in the decoupled feature space $\mathrm{\dot{E}_{i,m}}$:
(8) $\dot{\mathrm{E}}_{i,m}=\mathrm{W}_{1}\mathrm{E}_{i,m}+\mathrm{b}_{1},$
@@
where $\mathbf{W}_{1}\in\mathbb{R}^{d\times d_{\lambda}}$ and $\mathbf{b}_{1}\in\mathbb{R}^{d_{m}}$ represent trainable transformation matrix and bias vector, respectively.
@@
Considering that the behavioral information fully reflects the users’ preference, we filter the preference-related content features with the help of behavioral information guidance:
(9) $\ddot{\mathrm{E}}_{i,m}=f_{gate}^{m}(\mathrm{E}_{i,id},\dot{\mathrm{E}}_{i,m})=\mathrm{E}_{i,id}\odot\sigma(\mathrm{W}_{2}\dot{\mathrm{E}}_{i,m}+\mathrm{b}_{2}),$
@@
where $\mathbf{W}_{2}\in\mathbb{R}^{d_{\lambda}\times d_{\lambda}}$ and $\mathbf{b}_{2}\in\mathbb{R}^{d_{m}}$ are trainable parameters. Here, $\odot$ denotes element-wise multiplication, and $\sigma$ is the Tanh non-linear transformation.
Finally, the decoupled content features are recombined. Given that behavioral information reflects user preferences, we determine the weights of different decoupled content features based on the guidance provided by this behavioral information:
(10) $\overline{\mathrm{E}}_{i,m}=f_{merge}^{m}(\mathrm{E}_{i,id},\ddot{\mathrm{E}}_{i,m})=\mathrm{E}_{i,id}\odot\sigma(\mathrm{W}_{3}\ddot{\mathrm{E}}_{i,m}+\mathrm{b}_{3}),$
@@
where $\mathbf{W}_{3}\in\mathbb{R}^{d_{\lambda}\times d}$ and $\mathbf{b}_{3}\in\mathbb{R}^{d}$ are trainable parameters. Similarly, $\odot$ denotes element-wise multiplication, and $\sigma$ is the sigmoid non-linear transformation. It is important to note that although Equations 9 and 10 are formally similar, they achieve different effects.
@@
5. EXPERIMENTS
In this section, we conduct extensive experiments to evaluate the performance of the proposed adapter on three public datasets to answer following research questions.
  • RQ1: How effective is BeFA applied to existing multimodal recommender systems?
  • RQ2: How effective is BeFA in adapting content features?
  • RQ3: Why purifying modality information can achieve better recommendation performance?
  • RQ4: How do hyper-parameter settings impact the performance of BeFA?
5.1. Experimental Settings
5.1.1. Datasets
We conducted experiments on three publicly available datasets: (a) TMALL [^1] https://tianchi.aliyun.com/dataset/140281; (b) Microlens [^2] https://recsys.westlake.edu.cn/MicroLens-50k-Dataset/; and (c) H&M [^3] https://www.kaggle.com/datasets/odins0n/handm-dataset-128x128. We performed 10-core filtering on the raw data. The detailed information of the filtered data is presented in Table 1. For multimodal information, we utilized pre-trained CLIP (clip, ) and ImageBind (imagebind, ) to generate aligned multimodal features.
5.1.2. Comparative Model Evaluation
To verify the prevalence of feature encoder defects, we used ViT-B/32-based CLIP (clip, ) and ImageBind (imagebind, ) for our experiments. CLIP is a general-purpose feature encoder capable of handling a wide range of data types and application domains. It realizes cross-modality feature representations and exhibits highly efficient performance, making it widely adopted in practical applications and performing well across various tasks. ImageBind is the latest cross-modality feature encoder, which can handle a wider variety of modality information, and achieves a new SOTA performance in the emerging zero-shot recognition task for various modalities, even outperforming the previous specialized models trained specifically for these modalities.
To evaluate the effectiveness of our proposed adapter, we applied it to several representative multimodal recommendation models, including LATTICE (lattice, ), BM3 (bm3, ), FREEDOM (FREEDOM, ), and MGCN (mgcn, ) for comparison. Additionally, we compared our adapter with existing efficient parameter tuning methods such as LoRA (lora, ) and Soft-Prompt Tuning (softprompt, ).
Table 3. Performance Comparison with other Efficient Parameter Adaptation methods. The best result is in boldface and the second best is underlined. The t-tests validate the significance of performance improvements with p-value $\leq$ 0.05.
@@ 5.1.2. Comparative Model Evaluation
i) Multimodal Recommendation Models:
  • LATTICE (lattice, ): This method designs a new modality-aware structure learning layer that learns the item-item structure of each modality and aggregates multiple modalities to obtain latent item graphs. Based on the learned latent graphs, graph convolution is performed to explicitly inject higher-order item affinities into item representations.
  • BM3 (bm3, ): This method proposes a simple self-supervised learning framework that generates target views of users or items for comparison learning without negative samples.
  • FREEDOM (FREEDOM, ): This method devises a degree-sensitive edge pruning method to denoise user-item interaction graphs, rejecting potentially noisy edges with high probability when sampling the graph.
  • MGCN (mgcn, ): This method introduces a behavior-aware fuser to effectively model user preferences by dynamically learning the relative significance of various modality features. Additionally, it integrates the fuser with a self-supervised auxiliary task aimed at maximizing the mutual information between the fused multimodal features and behavioral features. This approach enables the concurrent capture of complementary and supplementary preference information.
ii) Efficient Parameter Tuning Methods
  • Low-Rank Adaptation (LoRA) (lora, ): This method aims to enhance the representation of a model while keeping the number of model parameters relatively small. It provides additional learning power by introducing lower rank matrices to linearly transform the hidden representation of the original model. This approach strikes a balance between parameter efficiency and model performance.
  • Soft Prompt Turning (softprompt, ): This method enhances model performance by fine-tuning small, task-specific embeddings (prompts) inserted into the input, thereby adapting pre-trained language models to new tasks without altering their original weights.
5.1.3. Evaluation Protocols
For a fair comparison, we follow the evaluation settings in (lattice, ; bm3, ) with the same 8:1:1 data splitting strategy for training, validation and testing. Besides, we follow the all-ranking protocol to evaluate the top-K recommendation performance and report the average metrics for all users in the test set: $\mathrm{R@K}$ and $\mathrm{N@K}$, which are abbreviations for $\mathrm{Recall@K}$ (recall, ) and $\mathrm{NDCG@k}$ (NDCG, ), respectively.
5.1.4. Implementation Details
We implement MMRec [^4] https://github.com/enoche/MMRec (mmrec, ) based on PyTorch, which is a unified public repository designed for multimodal recommendation methods. To ensure fair comparison, we employed the Adam optimizer to optimize all methods and referred to the best hyperparameter settings reported in the original baseline paper. For general settings, we initialized embeddings with Xavier initialization of dimension 64, set the regularization coefficient to $\lambda_{E}=10^{-4}$, and batch size to $B=2048$. Early stopping and total epochs are fixed at 10 and 1000, respectively. We selecte the best model with the highest $\mathrm{Recall@20}$ metric on the validation set and reported metrics on the test set accordingly.
5.2. Overall performance(RQ1)
5.2.1. Effectiveness
(a) Comparison of model performance before and after applying BeFA
Table 2 shows the performance comparison of BeFA applied to existing multimodal recommendation models on three datasets. From the table,we find several observations:
@@ 5.2.1. Effectiveness
  • Our adapter significantly improved recommendation performance on all three datasets. Specifically, our adapter has achieved an average improvement of 9.07% and 11.02% over the baseline in terms of $\mathrm{Recall@20}$ and $\mathrm{NDCG@20}$. This suggests that the content features extracted by the pre-trained encoder do have deficiencies that affect recommendation performance. Our processing of content features is very efficient and effectively reduces modality noise. This helps the multimodal recommendation models to better utilize the modality information and thus improve the recommendation performance.
  • Our adapter has significant performance improvements for all encoders. All pre-trained feature encoders are generally defective. We use CLIP and ImageBind as an example, and applying our adapter to the multimodal features extracted by these pre-trained feature encoders has led to a significant improvement in the performance of the recommender system. This suggests that the deficiencies are mainly due to the pre-training data, indicating that the content features extracted by the pre-trained encoders generally do not reflect the users’ preferences.
  • More powerful encoder can extract more sufficient content information. ImageBind has better overall preformance than CLIP when applied to recommendation models. This indicates that more advanced encoders are able to extract more sufficient content information. Thus better modelling of users and items to improve the performance of recommendations.
(b) Performance comparison with existing parameter tuning methods
Table 4. The training cost. #Param: number of tunable parameters, Time/E: averaged training time for one epoch, ’s’ means seconds. The results are collected on a GeForce RTX 4090 GPU.
@@
Figure 7. The effect of LoRA and SoftPrompt Turning on the adaptation of features. (a) is the original heatmap, (b) is the heatmap of visual features adapted by LoRA, (c) is the heatmap of visual features adapted by SoftPrompt Turning and (d) is the heatmap of visual features adapted by BeFA.
Table 3 shows the comparison between our adapter and existing efficient parameter adaptation methods. From the table,we find several observations:
  • The performance of our adapter is overall significantly better than existing efficient parameter tuning methods. This indicates that our adapter is better suited for the recommendation task than generalized efficient parameter tuning methods. It suggests that our adapter employs behavioral information guidance to more effectively obtain content features related to users’ preferences, thereby improving the effectiveness of recommendations.
  • Existing parameter tuning approaches are not generalized for recommendation tasks. The performance decreases in some cases instead, which indicates that these existing generalized parameter tuning methods are not necessarily suitable for the recommendation tasks. This highlights the importance of incorporating behavioral information guidance in recommendation tasks to improve model performance.
Figure 8. The gaps in the effectiveness of our adapter on different datasets. (a) is the original image. (b) is the heatmap of the visual features extracted by the pre-trained CLIP. (c) is the heatmap of visual features adapted by BeFA.
[…]
HM_Microlens
Figure 9. The Distribution of representations in different multimodal recommendation model in visual modalities before and after adaptation by BeFA.
5.2.2. Efficiency
We also analyze the efficiency of our adapter with other existing efficient parameter tuning methods on the BM3 model. The specific model parameter counts and training costs are shown in Table 4. Compared to the overall model training time, our training time increases by only about 15%, indicating that our method is not only effective but also efficient.
@@ 5.2.2. Efficiency
Although the design of our adapter is more complex compared to LoRA and Soft-Prompt Turning the increase in the number of parameters after using our adapter is still small compared to the overall parameters of the recommender system itself, ranging from 0.93% to 2.09%. In terms of training costs, the increase in overall training time after using our adapter is still within an acceptable range. The training time for the recommender system using our adapter is comparable to that of using LoRA and Soft-Prompt Tuning. This suggests that although our adapter introduces additional parameters, it does not impose a significant burden on the training process. The substantial improvement in performance of our adapter highlights its efficiency, trading a small number of parameters and training time gains for a significant improvement in recommendation accuracy.
5.3. Visualization Analysis(RQ2&RQ3)
5.3.1. Case Study
(a)Effectiveness with different models
To visually demonstrate the effect of our adapter on content features, we applied it to the proposed visualization attribution analysis. Specifically, before calculating the similarity between content features and behavioral features in the attribution analysis, we processed the content features with our adapter. We then computed similarity weights with the behavioral features and generated new heatmaps using these newly calculated weights with the corresponding masked images.
As shown in Figure 6, by analysing the heatmap before and after our adapter, we can clearly find that our adapter’s enhancement of content features is very obvious. Our adapter significantly improves the attention areas of features extracted by the pre-trained feature encoder, addressing issues of information drift and information omission. The adapted features focus more on the recommended products, with a marked decrease in attention to background elements unrelated to the recommendation. This reduction in noise enhances feature representation and consequently improves recommendation performance.
Additionally, the adapted features more comprehensively capture product details, significantly increasing the coverage of the area of interest for products. This comprehensive extraction of product details not only improves feature recognition but also enhances the recommender system’s ability to model product features. This means that the recommender system is able to more accurately capture the core features of the product, thus providing more relevant and accurate results when making recommendations. Overall,our adapter enables multimodal recommender systems to better utilize content features by using behavioral information to guide feature adaptation, thus significantly improving recommendation performance.
(b) Effectiveness with different encoders
@@ 5.3.1. Case Study
We utilize the content features extracted by two representative multimodal feature encoders, the ViT-32/B version of CLIP and ImageBind, and analyze the heatmaps of the content features within different downstream recommendation models after applying BeFA. Our analysis reveals that the content features extracted by both encoders exhibit issues of information drift and information omission. After applying BeFA, these content features are better focused on the recommended products themselves. This demonstrates that the deficiencies of pre-trained feature encoders are common in multimodal recommendation tasks, while our adapter effectively adapts to the content features extracted by different feature encoders, highlighting its general applicability.
(c) Comparison with existing adaptation methods
@@
We visualize and analyze the content features adapted by LoRA and Soft-Prompt Tuning, as shown in Figure 7. The results indicate that while LoRA and Soft-Prompt Tuning provide some optimization, the degree of adaptation is limited. The adapted content features still exhibit issues such as content error and content insufficiency, failing to focus accurately on the content of the product itself. These methods fail to adequately address the issues of information drift and information omission. This is the reason for the overall poor performance of existing adaptation methods in recommendation tasks. In contrast, the impact of BeFA on content features content features is significantly better than that of existing methods, which is why BeFA performs well in the recommendation task.
(d) Analysis in different scenarios
@@
We employed three datasets ranging from simple to complex scenarios. The H&M dataset comprises only product images with minimal interference, making it relatively clean. In contrast, the Microlens dataset comprises more complex images with more background interference. Our adapter demonstrated significant improvements across both datasets. Specifically, on the H&M dataset, it increases $\mathrm{Recall@20}$ and $\mathrm{NDCG@20}$ by 14.17% and 17.34%, respectively. Meanwhile, on the Microlens dataset, our adapter increases $\mathrm{Recall@20}$ and $\mathrm{NDCG@20}$ by 4.58% and 5.19%, respectively. These findings indicate the adapter’s effectiveness across varying levels of scene complexity. Moreover, as illustrated in Figure 8, even in complex scenarios, our adapter effectively removes irrelevant information and locates the recommended products as precisely as possible
5.3.2. Representation distribution analysis
To visually demonstrate the impact of BeFA on adjusting content features, we visualize the distribution of visual representations in the TMALL dataset. Figure 9 illustrates the comparison of our adapter’s application to feature distributions across various recommendation models. Specifically, we randomly sample 500 item representations from the TMALL dataset and map them to two-dimensional normalized vectors on the unit hypersphere $\mathcal{S}^{1}$ (i.e., a circle with a radius of 1) by using t-SNE (tsne, ). Subsequently, we employ Gaussian kernel density estimation(KDE) (KDE, ) to plot the distribution of these two-dimensional features.
Further examining the changes in the distribution of two dimensional features across different recommendation models before and after applying BeFA, we observe that the original feature distribution is more dispersed, forming multiple small dense regions. This distribution indicates that there is a large amount of noise in the visual feature extraction process, hindering the models’ ability to effectively distinguish key features of different products. In contrast, after applying BeFA, the feature distribution becomes more uniform and smooth, which indicates that our adapter has significant effects in eliminating noise. This improvement not only enhances feature recognition capability but also improves the model’s ability to distinguish different content features (pre, ).
Moreover, comparing the changes in feature distributions across different models, it can be found that BeFA has strong adaptability. The feature distributions of various multimodal recommendation models show significant improvements after applying BeFA. This indicates that our adapter is not only effective on a single model but also has strong generalization capabilities across different recommender systems. This adaptability highlights BeFA’s wide potential and value in practical applications.
5.4. Sensitivity Analysis(RQ4)
Figure 10. Performance comparison w.r.t. different size of the decoupling space relative to the original embedding space $\lambda$.
[…]
We investigate the size of the decoupling space relative to the original embedding space, as shown in Fig.10. The results show that the optimal size of the decoupling space is about four times the size of the original embedding space on some datasets. When the decoupling space dimension is small (e.g., 0.125x and 0.25x), it may create an information bottleneck, leading to poor performance across all datasets in both $\mathrm{Recall@20}$ and $\mathrm{NCGD@20}$. A smaller decoupling space cannot hold sufficient information, making it difficult for the model to accurately capture and differentiate features, thereby negatively impacting recommendation performance. Conversely, an excessively large decoupling space dimension may lead to dimensionality catastrophe, introducing noise and redundant features which complicates the model’s ability to effectively distinguish and learn features in a high-dimensional space. Additionally, its number of parameters and the computational cost will increase dramatically, bringing additional burden to the training process, which is impractical for practical application scenarios.
Furthermore, we find that there is a significant difference in the performance of Microlens, TMALL and H&M datasets under varying decoupling space sizes. For the Microlens dataset, both $\mathrm{Recall@20}$ and $\mathrm{NCGD@20}$ achieve the optimal performance when the decoupling space size is four times of the original embedding space. In contrast, in the TMALL dataset, although the overall trend is similar, the optimal point fluctuates slightly, suggesting that the size of the decoupling space may need to be adjusted on different datasets to obtain optimal results. The overall adapter performance fluctuates greatly with the change in the decoupling space size, highlighting the challenge of finding the optimal size. This is one of the limitations of our work, which could be addressed by considering the decoupling space size as a learnable parameter in future research. This approach could adaptively adjust the decoupling space size for different application scenarios to achieve the best performance.
6. Conclusion

### v2

@@ Introduction
ImageBind Avg Improve 15.54% 9.09% 17.99% 12.53% 5.18% 3.77% 4.83% 4.12% 15.24% 13.87% 19.56% 17.92%
Table 1: Performance Comparison on Different Recommender Models. The $t$-tests validate the significance of performance improvements with $p$-value $\leq$ 0.05.
Let $u\in\mathcal{U}$ and $i\in\mathcal{I}$ denote the user and item, respectively. The input behavioral features for user $u$ and item $i$ are represented as $\mathrm{E_{b}}\in\mathbb{R}^{d\times(|\mathcal{U}|+|\mathcal{I}|)}$, where $d$ is the embedding dimension. Each row represents the user embedding $\boldsymbol{h_{u}}$ and item embedding $\boldsymbol{h_{i}}$. They are initialised with ID information (He and McAuley 2016). Each item content feature is denoted as $\boldsymbol{e_{i,m}}\in\mathbb{R}^{d_{m}}$, where $d_{m}$ is the dimension of the features. $m\in\mathcal{M}$ represents the modality and $\mathcal{M}$ is the set of modalities. Multimedia recommendation aims to rank items for each user by predicting preference scores $\hat{y}_{u,i}$, which indicates the possibility of interaction between the user $u$ and the item $i$. The recommended process can be viewed as a function $f(\cdot)$. This process can be viewed simply as:
$\hat{y}_{u,i}={f}(\boldsymbol{h_{u}},\boldsymbol{h_{i}},\boldsymbol{e_{i,m}})$ (3)
Existing multimodal recommendation methods typically use pre-trained encoders to extract content features $\boldsymbol{e_{i,m}}$. Due to the deficiencies of the pre-trained encoder, a large deviation in the content features $\boldsymbol{e_{i,m}}$ leads to a bias in the prediction score $\hat{y}_{u,i}$, which affects the recommendation results. Therefore, we designed BeFA to adapt the content features to make them more suitable for recommendation. BeFA improves the recommendation results by adjusting $\boldsymbol{e_{i,m}}$, which can be viewed as a function $g(\cdot)$. The adjusted predicted preference scores is donated:
$\hat{y}_{u,i}={f}[\boldsymbol{h_{u}},\boldsymbol{h_{i}},g(\boldsymbol{h_{i}},\boldsymbol{e_{i,m}})]$ (4)
Behavior-driven Feature Adapter
@@
Due to the shortcomings of pre-trained feature encoders, their extracted content features contain amount of irrelevant and erroneous information. To better utilize modality information,we propose Behaviour-driven Feature Adapter(BeFA) for adapting content features. Firstly, we decouple the original item $i$ content features $\boldsymbol{e_{i,m}}$ in the decoupled feature space:
$\dot{\boldsymbol{e}}_{i,m}=\mathbf{W}_{1}\boldsymbol{e}_{i,m}+\mathbf{b}_{1},$ (5)
@@
where $\mathbf{}{W}_{1}\in\mathbf{}{R}^{d_{m}\times d_{a}}$ and $\mathbf{b}_{1}\in\mathbb{R}^{d_{a}}$ represent trainable transformation matrix and bias vector. $d_{a}$ is a hyper-parameter represents the dimension of the decoupled space.
@@
Considering that the behavioral information reflects the users’ preference (Yuan et al. 2023), we filter the preference-related content features with the guidance of behavioral information. The filter function $f_{gate}^{m}(\cdot)$ is represented as:
$\ddot{\boldsymbol{e}}_{i,m}=f_{gate}^{m}(\boldsymbol{b}_{i},\dot{\boldsymbol{e}}_{i,m})=\boldsymbol{b}_{i}\odot\tanh(\mathbf{W}_{2}\dot{\boldsymbol{e}}_{i,m}+\mathbf{b}_{2}),$ (6)
@@
where $\mathbf{W}_{2}\in\mathbb{R}^{d_{a}\times d_{a}}$ and $\mathbf{b}_{2}\in\mathbb{R}^{d_{a}}$ are trainable parameters, $\boldsymbol{b_{i}}$ denotes the behavioral information of item i. Here $\tanh(\cdot)$ is the Tanh non-linear transformation.
Finally, the decoupled content features are selectively recombined. By combining different decoupled features with behavioral guidance, we aim to enhance the system’s ability to capture nuanced and contextually relevant aspects of the content. The combination function $f_{merge}^{m}$ is represented as:
$\overline{\boldsymbol{e}}_{i,m}=f_{merge}^{m}(\boldsymbol{b}_{i},\ddot{\boldsymbol{e}}_{i,m})=\boldsymbol{b}_{i}\odot\sigma(\mathbf{W}_{3}\ddot{\boldsymbol{e}}_{i,m}+\mathbf{b}_{3}),$ (7)
@@
where $\mathbf{W}_{3}\in\mathbb{R}^{d_{a}\times d_{m}}$ and $\mathbf{b}_{3}\in\mathbb{R}^{d_{m}}$ are trainable parameters. $\sigma(\cdot)$ is the sigmoid non-linear transformation. It is important to note that although Equations 6 and 7 are formally similar, they achieve different effects.
@@
Experiments
Experimental Settings
Datasets
We conducted experiments on three publicly available datasets: (a) TMALL [^1] https://tianchi.aliyun.com/dataset/140281; (b) Microlens [^2] https://recsys.westlake.edu.cn/MicroLens-50k-Dataset/; and (c) H&M [^3] https://www.kaggle.com/datasets/odins0n/handm-dataset-128x128. The detailed information is presented in the Appendix. For multimodal information, we utilized pre-trained CLIP (Radford et al. 2021) and ImageBind (Girdhar et al. 2023) to extract aligned multimodal features.
Comparative Model Evaluation
To verify the prevalence of feature encoder defects, we used ViT-B/32-based CLIP and ImageBind for our experiments. CLIP and ImageBind are advanced cross-modality feature encoders. CLIP known for its broad applicability and efficiency, and ImageBind achieving SOTA performance in zero-shot recognition across various modalities. To evaluate the effectiveness of BeFA, we applied it to several representative multimodal recommendation models, including LATTICE (Zhang et al. 2021), BM3 (Zhou et al. 2023c), FREEDOM (Zhou and Shen 2023), and MGCN (Yu et al. 2023). Additionally, we compared BeFA with existing efficient parameter tuning methods, including LoRA (Hu et al. 2021) and Soft-Prompt Tuning (Lester, Al-Rfou, and Constant 2021).
Evaluation Protocols & Implementation
Due to the length of the article, the details are shown in the Appendix.
Overall performance
Figure 5: Visualization analysis of the effect of adapter on feature purification. The shade of the colour represents the amount of attention weight.
@@
Table 2: Performance Comparison with other Efficient Parameter Adaptation methods. The best result is in boldface and the second best is underlined. The $t$-tests validate the significance of performance improvements with $p$-value $\leq$ 0.05.
@@
The effectiveness of BeFA. Table 1 demonstrates that our adapter enhances recommendation performance across all three datasets. Specifically, our adapter has achieved average improvements of 9.07% and 11.02% over the baseline in terms of Recall@20 and NDCG@20. This suggests that the pre-extracted content features do have deficiencies that negatively impact recommendation performance. Our processing of content features effectively reduces modality noise. This helps the multimodal recommendation models to better utilize the modality information and thus improve the recommendation performance. Furthermore, Table 2 shows the performance of our adapter is better than existing efficient parameter tuning methods, highlighting its superior suitability for recommendation tasks. This suggests that our adapter uses behavioral information to better capture user-preferred content features, thus improving the performance.
The generalization of BeFA. From Table 1, we observe that BeFA improves performance across various multimodal recommendation models, highlighting its effectiveness as a generic plugin for all multimodal recommendations. Existing multimedia recommendation methods ignore this problem, thus leading to suboptimal results. Enhancing content features by BeFA, can make the content features better reflect the users’ preference information. Thus make the existing methods more accurately model the users’ preference. Meanwhile, BeFA enhances performance for all types of encoders, indicating thar pre-trained encoders are generally defective. It can also be found that although more advanced feature encoder (ImageBind) can be used to extract information more comprehensively. The extracted features still contain a large amount of irrelevant information, resulting in generally do not accurately reflect the user’s preference.
@@
Table 3: The training cost. #Param: number of tunable parameters, Time/E: averaged training time for one epoch.
@@
The efficiency of BeFA. From Table 3, we observe that our adapter is more complex compared to LoRA and Soft-Prompt Tuning. However, the increase in the number of parameters after using our adapter remains small, ranging from 0.93% to 2.09% of the overall recommender system’s parameters. The increase in overall training time is only about 15%, which is comparable with other existing methods. This suggests that although our adapter is more complex and introduces additional parameters, it does not impose a burden on the training process.The substantial improvement in performance of our adapter highlights its efficiency. It trades a small increase in parameters and training time for a improvement in recommendation performance.
Visualization Analysis
Figure 6: Comparison between different parameter-efficient adaptation methods.
BeFA guides content features to more accurately reflect user preferences. To visually demonstrate the effect of our adapter on content features, we applied the purified features to the proposed visualization attribution analysis. As shown in Figure 5, our adapter addresses issues of information drift and information omission. The adapted features focus more on the recommended items, with a marked decrease in attention to background elements unrelated to the recommendation. Additionally, the adapted features more comprehensively capture item details. This comprehensive extraction of item details improves feature recognition, which enhances the recommender system’s ability to model item features. Thus provides more relevant and accurate results when making recommendations. We also visualized the effect of BeFA on the distribution of visual representations. We find that the adapted features more accurately reflect the detailed information of the items, and the discriminability of the features is clearly improved. This is reflected in the distribution as a noticeable improvement in homogeneity (Wang et al. 2022).
@@
BeFA is generalized with different encoders. We utilize the content features extracted by two representative multimodal feature encoders, the ViT-32/B version of CLIP and ImageBind, and analyze the heatmaps of the content features within different downstream recommendation models after applying BeFA. Our analysis reveals that the content features extracted by both encoders exhibit issues of information drift and information omission. After applying BeFA, these content features are better reflect the recommended items themselves. This demonstrates that the deficiencies of pre-trained feature encoders are common in multimodal recommendation tasks, while our adapter effectively adapts to the content features extracted by different feature encoders, highlighting its general applicability.
Figure 7: The effectiveness of BeFA on different scenarios. The top image is the heatmap and the bottom one is the heatmap after applying BeFA.
@@
BeFA outperformance than other parameter tuning methods. We visualize and analyze the content features adapted by LoRA and Soft-Prompt Tuning, as shown in Figure 6. The results indicate that although LoRA and Soft-Prompt Tuning offer some optimization, the degree of adaptation is limited. These methods do not sufficiently address issues of information drift and omission, failing to focus accurately on the content of the item itself. This explains why existing adaptation methods perform poorly in recommendation tasks. In contrast, BeFA has a greater impact on content characterisation than existing methods. The quality of its adjusted content features is obviously higher. It can better model the items and thus improve the performance.
@@
BeFA performs well across various scenarios. We employed three datasets ranging from simple to complex scenarios. The H&M dataset comprises only item images with minimal interference, making it relatively clean. In contrast, the Microlens dataset comprises more complex images with more background interference. Our adapter demonstrated improvements across both datasets. Specifically, on the H&M dataset, it increases Recall@20 and NDCG@20 by 14.17% and 17.34%. Meanwhile, on the Microlens dataset, our adapter increases Recall@20 and NDCG@20 by 4.58% and 5.19%. BeFA reduces the interference between content features and filters and retains the most relevant features for the recommendation task, as illustrated in Figure 7. As a result, it achieves better recommendation results even in various complex situations.
Hyperparameter Analysis
Due to the length of the article, specific analyses are provided in the Appendix.
Conclusion